OpenStudy (anonymous):

How would I convert (square root of 10)^3 to exponential form?

OpenStudy (campbell_st):

so for roots you need fractional indices... \[\sqrt[a]{x} = x^{\frac{1}{a}}\] then use the power of a power law, where you multiply the powers \[(x^a)^b = x^{a \times b}\] hope it helps

OpenStudy (anonymous):

Thank you so much! It helps loads!